# Differential Test Performance and Peer Effects

### v1

@@ 4.1 Identification Conditions
Let Assumptions 1, 2, 3, 4, 5, 6 and 7 be satisfied. Let $m_{n}\left(\theta,\gamma\right)$ be defined as in (29). Then for all admissible $\gamma\in\Gamma,$ $E\left[m_{n}\left(\theta_{0},\gamma\right)\left|\mathcal{Z}_{n}\right.\right]=0$ and $E\left[m_{n}\left(\theta,\gamma\right)\left|\mathcal{Z}_{n}\right.\right]\neq 0$ for $\theta\neq\theta_{0}$ a.s.
4.2 Consistency of Parameter Estimators
@@
Formulating the moment conditions using transformed residuals as in (29) requires consistent estimates of the variance parameters $\gamma$. These can be obtained by first obtaining consistent but inefficient parameter estimates. To do so, we set all elements of $\gamma=(\gamma_{1},...,\gamma_{J})^{\prime}$ to one. Lemma 4.2 implies that the parameters $\theta=(\rho,f_{1},\delta^{\prime})^{\prime}$ are still identified by the moment conditions in (29).
@@
First step estimators $(\tilde{\rho},\tilde{f}_{1},\tilde{\delta}^{\prime})^{\prime}$ are defined in (19) and (22). [^13] Lee (2007a) considers first step and efficient GMM estimators using linear and quadratic moments. To formalize the identification result for these non-efficient first step estimators define the moment vector
@@
Lemma 4.3.
Let Assumptions 1-6, and Assumption 7(i),(ii) be satisfied. Let $m_{\epsilon}(\theta)$ be defined as in (31). Assume further that $n^{-1}\left|E\left[z^{\prime}Q_{X}y_{2}|\mathcal{Z}_{n}\right]\right|>K_{y}>0$ a.s . Then $E\left[m_{\epsilon}\left(\theta_{0}\right)\left|\mathcal{Z}_{n}\right.\right]=0$ and $E\left[m_{\epsilon}\left(\theta\right)\left|\mathcal{Z}_{n}\right.\right]\neq 0$ a.s. if $\theta\neq\theta_{0}$ for all $\theta\in\Theta$
@@
Let $i=n_{1}+\ldots n_{c-1}+r$ with $1\leq r\leq n_{c}$. Then, with some abuse of notation, denote the $i$-th diagonal element of $\Omega(\gamma)=\operatorname{diag}_{c=1}^{C}\left(\gamma_{\tau_{c}}^{2}I_{c}\right)$ as $\gamma_{(i)}^{2}=\gamma_{\tau_{c}}^{2}$, noting that $\tau_{c}$ varies with $i$. Let the linear and quadratic moment vectors $m_{n}^{(l)}(\theta,\gamma)$ and $m_{n}^{(q)}(\theta,\gamma)$ be as defined in (29). Let $h_{i}=\left[h_{i1},...,h_{iq}\right]$ denote the $i$-th row vector of $H$, and let $a_{ij}$ denote the the $(i,j)$-th element of matrix $A$. Then note that with $u^{+}(\theta_{0},\gamma)=\Omega(\gamma)^{-1}(u_{1}-f_{1,0}u_{2})$, the variance of $m_{n}^{(l)}(\theta_{0},\gamma)$ is
@@
$\frac{1}{n}E\left[\operatorname{tr}\left(\Omega_{0}\Omega^{-1}A\Omega_{0}\Omega^{-1}A\right)\right]=\frac{1}{n}E\sum_{i=1}^{n}\sum_{j=1}^{n}\left[\frac{\gamma_{(i),0}^{2}}{\gamma_{(i)}^{2}}\frac{\gamma_{(j),0}^{2}}{\gamma_{(j)}^{2}}a_{ij}^{2}\right].$ (33)
@@
(i) Let $C\left(\theta,\gamma\right)$ be an $n\times n$ matrix of the form $D$, $DP$, $DAD$, $DADP$, or $PDADP$, where $D$ is an $n\times n$ positive diagonal matrix with elements which are uniformly bounded and measurable w.r.t. $\mathcal{Z}_{n}$ and where $P=\operatorname{diag}_{c}\left(p_{c}I_{c}^{\ast}+q_{c}J_{c}^{\ast}\right)$ and where $p_{c}$ and $q_{c}$ are continuously differentiable functions of the parameters $\theta$ and $\gamma$ as well as of variables generating $\mathcal{Z}_{n}$, and $I_{c}^{\ast}$ and $J_{c}^{\ast}$ are defined in Appendix B. Then
@@
Theorem 4.1.
Let Assumptions 1- 9 be satisfied. Assume that $\bar{\gamma}_{n}\stackrel{{\scriptstyle p}}{{\rightarrow}}\gamma$ for some sequence $\bar{\gamma}_{n}$ and some $\gamma\in\Gamma.$ Then,

### v2

@@ 5 Identification
Let Assumptions 3, 4, 5, 6, 7 and 8 be satisfied. Let $m_{n}\left(\theta,\gamma\right)$ be defined as in (19). Then for all admissible $\gamma\in\Gamma,$ $E\left[m_{n}\left(\theta_{0},\gamma\right)\left|\mathcal{Z}_{n}\right.\right]=0$ and $E\left[m_{n}\left(\theta,\gamma\right)\left|\mathcal{Z}_{n}\right.\right]\neq 0$ for $\theta\neq\theta_{0}$ a.s.
We report results that establish the consistency and asymptotic normality of our estimators in an online appendix where we also discuss consistent estimation of $\gamma$ and standard errors.
6 Conclusions
Appendix
Appendix A Proofs
Proof of Theorem 2.1.
By the assumption on the boundedness of the $\left(4+\eta\right)$-th moments all considered expectations and conditional expectations exist. Since $u_{it}=y_{it}^{*}-E\left[y_{it}^{*}|\mathcal{F}_{n,i,t}\right]$ we have $E\left[u_{it}|\mathcal{F}_{n,i,t}\right]=0$ by construction. Observing that $E\left[y_{it}^{*}|\mathcal{F}_{n,i,t}\right]=\kappa_{i}$ and that $\kappa_{i}$ is $\mathcal{Z}_{n}^{*}$-measurable it follows that conditional on $\mathcal{Z}_{n}^{*}$ knowledge of $y_{it}^{*}$ is equivalent to knowledge of $u_{it}$. To see this let $u_{-i,t}=\left(u_{1t},..,u_{i-1,t},u_{i+1,t},...,u_{n,t}\right)^{\prime}$ and use the notation $u_{-t}$ for the unobserved components related to the other test. Define $U=\left(S,\alpha,\varphi,\zeta^{*},u_{-t},u_{-i,t},\kappa_{i}\right)$ assuming for simplicity of notations that vectors are row vectors, and let $V=\left(S,\alpha,\varphi,\zeta^{*},y_{-t}^{*},y_{-i,t}^{*},\kappa_{i}\right)$. Then $V=g(U)$ where $g(.)$ is Borel measurable and thus $\sigma(V)\subseteq\sigma(U)$. Analogously $U=h(V)$ where $h(.)$ is Borel measurable and thus $\sigma(U)\subseteq\sigma(V)$. Consequently $\sigma(U)=\sigma(V)$. Since $\kappa_{i}$ is measurable w.r.t. $\sigma(S,\alpha,\varphi,\zeta^{*})$ it follows that
$\sigma(S,\alpha,\varphi,\zeta^{*},u_{-t},u_{-i,t},\kappa_{i})=\sigma(S,\alpha,\varphi,\zeta^{*},u_{-t},u_{-i,t})$
and $\sigma(S,\alpha,\varphi,\zeta^{*},y_{-t}^{*},y_{-i,t}^{*},\kappa_{i})=\sigma(S,\alpha,\varphi,\zeta^{*},y_{-t}^{*},y_{-i,t}^{*}),$ which shows that
$\mathcal{B}_{n,i,t}=\sigma\left(S,\alpha,\varphi,\zeta^{*},u_{-t},u_{-i,t}\right)=\sigma\left(S,\alpha,\varphi,\zeta^{*},y_{-t}^{*},y_{-i,t}^{*}\right)=\mathcal{F}_{n,i,t}.$
Next observe that the $u_{jt}$ are $\mathcal{F}_{n,i,t}$-measurable for $i\neq j$, and hence by iterated expectations $\operatorname{Cov}\left(u_{it},u_{jt}|\mathcal{F}_{n,i,t}\right)=E\left[u_{jt}E\left(u_{it}|\mathcal{F}_{n,i,t}\right)\right]=0$. Observe further that for $s\neq t$ the $u_{js}$ are $\mathcal{F}_{n,i,t}$-measurable for all $i$, and hence by iterated expectations $\operatorname{Cov}\left(u_{it},u_{js}|\mathcal{F}_{n,i,t}\right)=E\left[u_{js}E\left[u_{it}|\mathcal{F}_{n,i,t}\right]\right]=0$. The corresponding claims regarding the mean and covariances of the $u_{it}$ conditional on $\mathcal{G}_{n,i,t}\subseteq\mathcal{F}_{n,i,t}$ follow immediately by iterated expectations. This proves the first part of the theorem.
E$\left[y_{it}^{*}-\kappa_{i}|\mathcal{G}_{n,i,t}\right]$ $=$ $E\left[E\left[y_{it}^{*}|\mathcal{F}_{n,i,t}\right]|\mathcal{G}_{n,i,t}\right]-E\left[\kappa_{i}|\mathcal{G}_{n,i,t}\right]=E\left[\kappa_{i}|\mathcal{G}_{n,i,t}\right]-E\left[\kappa_{i}|\mathcal{G}_{n,i,t}\right]=0.$
Thus by the almost sure uniqueness of conditional expectations, Assumption 2 cannot hold if $E\left[y_{it}^{*}-\kappa_{i}|\mathcal{G}_{n,i,t}\right]\neq 0$. ∎
Note that Lemma 5.1 and Corollary 5.1 can be viewed as special cases of Theorem 5.1. We therefore only give an explicit proof of Theorem 5.1 here. Proofs of the additional results are given in the online appendix. [^14] Equations labeled 32 and higher, as well as Appendices B and higher can be found in the online appendix.
Proof of Theorem 5.1.
In light of (13) it follows that,
$y_{1}-f_{1}y_{2}-\text{\text@underline{X}}\delta=\mu^{\ast}(f_{1,0}-f_{1})+\underline{X}\left(\delta(f_{1},\rho_{0},\beta_{0})-\delta\right)+(I+\rho_{0}M)(u_{1}-f_{1}u_{2}).$ (27)
Let $\mathcal{M}(\rho)=(I+\rho M)^{-1}(I+\rho_{0}M)$. Substitution of (27) into (18) yields
$u^{+}(\theta,\gamma)$ $=$ $\Omega(\gamma)^{-1/2}(I+\rho M)^{-1}$
recalling that $V(\rho,\gamma)=\Omega(\gamma)^{-1/2}(I+\rho M)^{-1}$
We first analyze the linear moment conditions. Note that
Let
Observe that
where the r.h.s. is seen to hold in light of (33) and (35), and since $H$ is $\mathcal{Z}_{n}$-measurable and thus
The assumptions imply that $K_{n}$ has full rank a.s. To see this, note that in light of (73), (74), Lemma E.1(iv) in online Appendix E implies that
$V(\rho,\gamma)=\operatorname{diag}_{c=1}^{C}\left(\frac{1}{\gamma_{\tau_{c}}}\frac{n_{c}-1-\rho}{n_{c}-1}I_{c}^{\ast}+\frac{1}{\gamma_{\tau_{c}}}\left(1+\rho\right)J_{c}^{\ast}\right),$ (39)
Under Assumptions 5 and 7, there exists some constants $c_{e}$ and $C_{e}$ such that $0<c_{e}<\frac{1}{\gamma_{\tau_{c}}}\frac{n_{c}-1-\rho}{n_{c}-1}<C_{e}<\infty$ and $0<c_{e}<\frac{1}{\gamma_{\tau_{c}}}\left(1+\rho\right)<C_{e}<\infty$. In light of Lemma E.1(iii), the eigenvalues of $V(\rho,\gamma)$ are uniformly bounded below by $c_{e}>0$. By Assumption 8 we have $\lambda_{min}\left(n^{-1}\underline{X}^{\prime}\underline{X}\right)\geq\xi_{X}>0$ and thus $\lambda_{min}\left(n^{-1}\underline{X}^{\prime}V(\rho,\gamma)\underline{X}\right)\geq c_{e}\xi_{X}>0$ uniformly for all $\gamma$, $\rho$, and $n$.
Observe furthermore that $E\left[(V^{1/2}z)^{\prime}Q_{V^{1/2}X}(V^{1/2}y_{2})\left|\mathcal{Z}_{n}\right.\right]=E\left[(V^{1/2}z)^{\prime}Q_{V^{1/2}X}(V^{1/2}\mu^{*})\left|\mathcal{Z}_{n}\right.\right]$ in light of (13) and since $Q_{V^{1/2}X}$ is orthogonal to $V^{1/2}\underline{X}$ and $E\left[u_{i2}|\mathcal{Z}_{n}\right]=0$ by Theorem 2.1. Thus
$\inf_{\gamma\in\Gamma,\rho\in[-K_{\rho},K_{\rho}]}\left|\operatorname{det}\left(E\left[K_{n}(\rho,\gamma)\left|\mathcal{Z}_{n}\right.\right]\right)\right|\geq$ (44)
$\left(\xi_{X}c_{e}\right)\left(\inf_{\gamma\in\Gamma,\rho\in[-K_{\rho},K_{\rho}]}n^{-1}\left|E\left[(V^{1/2}z)^{\prime}Q_{V^{1/2}X}(V^{1/2}y_{2})\left|\mathcal{Z}_{n}\right.\right]\right|\right)$
$\geq\left(\xi_{X}c_{e}\right)^{p_{x}}K_{y}>0$
Once $f_{1,0}$ and $\delta_{0}$ are identified, $\rho$ can be identified from the quadratic moment condition. To see this, note that when evaluated at $\delta_{0}$ and $f_{1,0},$ $u^{+}(\theta,\gamma)$ in Equation (A) becomes
$=\Omega(\gamma)^{-1/2}\mathcal{M}(\rho)(u_{1}-f_{1,0}u_{2}).$
Then,
[…]
Online Appendix for “Differential Test Performance and Peer Effects” by Guido M. Kuersteiner, Ingmar R. Prucha and Ying Zeng
Appendix B Assumptions and Results for $u_{it}$
We impose the following formal assumption about the variance of the error structure.
Assumption 9.
Let $u_{it}=y_{it}^{*}-\kappa_{i}f_{t}$ and let $i=n_{1}+\ldots+n_{c-1}+r$ with $1\leq r\leq n_{c}$, i.e., $i$ refers to student $r$ in class c. Let $u_{-i,t}=\left(u_{1,t},...,u_{i-1,t},u_{i+1,t},...,u_{n,t}\right)$. Define the filtration $\mathcal{B}_{n,i,t}=\sigma\left(S,\alpha,\varphi,X,z,\zeta^{*},u_{-t},u_{-i,t}\right)\text{\,for $t=1,2$}$ where $u_{-t}$ represents information about the other test. Then the conditional variance of $u_{it}$ is given by $E\left[u_{it}^{2}|\mathcal{B}_{n,i,t}\right]=\sigma_{t}^{2}\rho_{\tau_{c}}^{2},$where $\tau_{c}$ is $\mathcal{Z}_{n}$-measurable and $\tau_{c}\in\{1,\ldots,J\}$, and $c_{u}\leq\sigma_{t}^{2},\rho_{j}^{2}\leq C_{u}$ for some finite positive bounds.
Assumption 9 implies that for $t=1,2$, $\Sigma_{t}\equiv E\left[u_{t}u_{t}^{\prime}|\mathcal{Z}_{n}^{\ast}\right]=\sigma_{t}^{2}\operatorname{diag}_{c=1}^{C}\left(\rho_{\tau_{c}}^{2}I_{c}\right)$. Observing that $f_{t}$ is a fixed constant, conditional on $S,\zeta^{*},\alpha,X$ and $z$, knowledge of $y_{it}^{*}$ is equivalent to knowledge of $u_{it}$. Thus we have the following equivalence of information sets [^15] See the proof of Lemma B.1 for more detail.
=$\mathcal{F}{}_{n,i,t}=\mathcal{B}_{n,i,t}\,\text{for}\,t=1,2.$
Assumption 9 imposes a similar factor structure on conditional variances as does Assumption 2 for conditional means. By Theorem 2.1 the errors $u_{it}$ are mutually uncorrelated over tests as well as cross-sectionally. Consequently the variance covariance matrix of $u_{1}-f_{1}u_{2}$ is given by,
Define $\gamma_{j}^{2}=\left(\sigma_{1}^{2}+f_{1}^{2}\sigma_{2}^{2}\right)\rho_{j}^{2}$, for $j=1,...,J$ and $\gamma=\left(\gamma_{1},...,\gamma_{J}\right)^{\prime}$. Then $\Omega(\gamma)=\operatorname{diag}_{c=1}^{C}\left(\gamma_{\tau_{c}}^{2}I_{c}\right)$ and the conditions about $\Omega(\gamma)$ and $\gamma$ in Assumption 5 hold.
Lemma B.1.
Let Assumption 4 hold. Then, $u_{it}$ defined in Theorem 2.1 is a martingale difference sequence w.r.t to the filtration $\mathcal{A}_{n,\nu}$ with $\mathcal{A}_{n,0}=\mathcal{Z}_{n}^{*}=\sigma\left(S,\alpha,\varphi,X,z,\zeta^{*}\right)$ and
$\mathcal{A}_{n,\left(t-1\right)n+i-1}=\sigma\left(S,\alpha,\varphi,X,z,\zeta^{*},\left\{u_{j,t-1}\right\}_{j=1}^{n},\left\{u_{jt}\right\}_{j=1}^{i-1}\right)\text{\,for $t=1,2;i=1,\ldots,n$}$ (45)
Proof.
Appendix C Consistency of Parameter Estimators
@@
Formulating the moment conditions using transformed residuals as in (19) requires consistent estimates of the variance parameters $\gamma$. These can be obtained by first obtaining consistent but inefficient parameter estimates. To do so, we set all elements of $\gamma=(\gamma_{1},...,\gamma_{J})^{\prime}$ to one. Theorem 5.1 implies that the parameters $\theta=(\rho,f_{1},\delta^{\prime})^{\prime}$ are still identified by the moment conditions in (19).
@@
First step estimators $(\tilde{\rho},\tilde{f}_{1},\tilde{\delta}^{\prime})^{\prime}$ are defined in (14), (15) and (16). [^16] Lee (2007a) considers first step and efficient GMM estimators using linear and quadratic moments. To formalize the identification result for these non-efficient first step estimators define the moment vector
@@
Lemma C.1.
Let Assumptions 3-7, and Assumption 8(i),(ii) be satisfied. Let $m_{\epsilon}(\theta)$ be defined as in (47). Assume further that $n^{-1}\left|E\left[z^{\prime}Q_{X}y_{2}|\mathcal{Z}_{n}\right]\right|>K_{y}>0$ a.s . Then $E\left[m_{\epsilon}\left(\theta_{0}\right)\left|\mathcal{Z}_{n}\right.\right]=0$ and $E\left[m_{\epsilon}\left(\theta\right)\left|\mathcal{Z}_{n}\right.\right]\neq 0$ a.s. if $\theta\neq\theta_{0}$ for all $\theta\in\Theta$.
Proof of Lemma C.1.
Observe that utilizing (26) we have
Observe that $E\left[\tilde{K}_{n}\right]$ has full rank, as is readily seen from the proof of Theorem 5.1 in light of (44) with $V=I$. This implies that the only solution to (48) is $\delta=\delta_{0}$ and $f_{1}=f_{1,0}$. With $\mathcal{M}(\rho)=(I+\rho M)^{-1}(I+\rho_{0}M)$, it now follows that
such that the quadratic moment condition is
By Lemma E.2 it follows that the only solution to (49) is $\rho=\rho_{0}.$
@@
Let $i=n_{1}+\ldots n_{c-1}+r$ with $1\leq r\leq n_{c}$. Then, with some abuse of notation, denote the $i$-th diagonal element of $\Omega(\gamma)=\operatorname{diag}_{c=1}^{C}\left(\gamma_{\tau_{c}}^{2}I_{c}\right)$ as $\gamma_{(i)}^{2}=\gamma_{\tau_{c}}^{2}$, noting that $\tau_{c}$ varies with $i$. Let the linear and quadratic moment vectors $m_{n}^{(l)}(\theta,\gamma)$ and $m_{n}^{(q)}(\theta,\gamma)$ be as defined in (19). Let $h_{i}=\left[h_{i1},...,h_{iq}\right]$ denote the $i$-th row vector of $H$, and let $a_{ij}$ denote the the $(i,j)$-th element of matrix $A$. Then note that with $u^{+}(\theta_{0},\gamma)=\Omega(\gamma)^{-1/2}(u_{1}-f_{1,0}u_{2})$, the variance of $m_{n}^{(l)}(\theta_{0},\gamma)$ is
@@
$\frac{2}{n}E\left[\operatorname{tr}\left(\Omega_{0}\Omega^{-1}A\Omega_{0}\Omega^{-1}A\right)\right]=\frac{2}{n}E\sum_{i=1}^{n}\sum_{j=1}^{n}\left[\frac{\gamma_{(i),0}^{2}}{\gamma_{(i)}^{2}}\frac{\gamma_{(j),0}^{2}}{\gamma_{(j)}^{2}}a_{ij}^{2}\right].$ (51)
@@
(i) Let $C\left(\theta,\gamma\right)$ be an $n\times n$ matrix of the form $D$, $DP$, $DAD$, $DADP$, or $PDADP$, where $D$ is an $n\times n$ positive diagonal matrix with elements which are uniformly bounded and measurable w.r.t. $\mathcal{Z}_{n}$ and where $P=\operatorname{diag}_{c}\left(p_{c}I_{c}^{\ast}+q_{c}J_{c}^{\ast}\right)$ and where $p_{c}$ and $q_{c}$ are continuously differentiable functions of the parameters $\theta$ and $\gamma$ as well as of variables generating $\mathcal{Z}_{n}$, and $I_{c}^{\ast}$ and $J_{c}^{\ast}$ are defined in Appendix E. Then
@@
Theorem C.1.
Let Assumptions 3- 11 be satisfied. Assume that $\bar{\gamma}_{n}\stackrel{{\scriptstyle p}}{{\rightarrow}}\gamma$ for some sequence $\bar{\gamma}_{n}$ and some $\gamma\in\Gamma.$ Then,